OpenStudy (anonymous):

Construct an equation and then use it to solve the problem. (a) Pressure P is directly proportional to temperature T, and P is 25 lb per square inch when T is 75 degrees Fahrenheit. What is the pressure when the temperature is 91 degrees Fahrenheit? Round answer to 3 decimal places.

OpenStudy (ybarrap):

P = kT kT = P k = P/T = 25/75, since P = 25 lb/in^2 and T = 75 = 1/3 P(91) = k(91) = (1/3)(91) = 91/3 Ans. P(91) is about 30.333